# DepsRAG: Towards Managing Software Dependencies using Large Language Models

### v1

@@ Abstract.
Managing software dependencies is a crucial maintenance task in software development and is becoming a rapidly growing research field, especially in light of the significant increase in software supply chain attacks. Specialized expertise and substantial developer effort are required to fully comprehend dependencies and reveal hidden properties about the dependencies (e.g., number of dependencies, dependency chains, depth of dependencies) .
Recent advancements in Large Language Models (LLMs) allow the retrieval of information from various data sources for response generation, thus providing a new opportunity for uniquely managing software dependencies. To highlight the potential of this technology, we present DepsRAG, a proof-of-concept Retrieval Augmented Generation (RAG) approach that constructs direct and transitive dependencies of software packages as a Knowledge Graph (KG) in four popular software ecosystems. DepsRAG can answer user questions about software dependencies by automatically generating necessary queries to retrieve information from the KG, and then augmenting the input of LLMs with the retrieved information. DepsRAG can also perform Web search to answer questions that the LLM cannot directly answer via the KG. We identify tangible benefits that DepsRAG can offer and discuss its limitations. [^1] DepsRAG demo and implementation are available here: https://shorturl.at/npyIZ.
copyright: none
@@ 1. Introduction
Software development increasingly relies on code reuse, which can be accomplished through software packages. Utilizing third-party packages to build software can significantly improve quality and productivity (Lim, 1994; Mohagheghi et al., 2004). These packages, dependencies, and their maintainers have formed large software ecosystems (Sonatype, 2021). Managing dependencies between packages is an emerging and important problem (Artho et al., 2012; Bogart et al., 2016; Decan et al., 2019), especially as open-source and third-party components have become widely adopted in software development. This new trend highlights the importance of software supply chain security, which is becoming more critical over time.
Today, developers are expected to manage their software dependencies (Kula et al., 2018). However, in a recent survey, developers reported that they are not confident in their current dependency management practices (Tidelift, [n. d.]). While developers often have full control over which dependency packages and versions to use, they have no control over the dependencies of the used packages. Such transitive dependencies, which often amount to more than direct dependencies, may have various vulnerabilities that put software projects at risk. To mitigate these risks, we developed DepsRAG to support practitioners and software developers understand the dependencies in their software supply chain. DepsRAG facilitates constructing and analyzing software dependencies for four popular software ecosystems. It also answers key questions about the dependency graph, including the number of dependencies, the depth of the dependency graph, and the number of paths/chains. The use of LLM enables DepsRAG to autonomously generate queries for KGs, unlike prior methods demanding hardcoded or user-generated queries (Dann et al., 2023; Litzenberger et al., 2023; Yamaguchi et al., 2014), thereby enhancing scalability. This retrieved information enhances the user’s question, referred to as Retrieval Augmented Generation (RAG), enabling accurate LLM responses. DepsRAG works in two phases. In the first phase, the user provides the name and ecosystem of the project. Then it transitively constructs the dependencies as a KG. In the second phase, the user can chat with DepsRAG to answer questions about the dependencies. DepsRAG is equipped with several tools to facilitate its functionality to access the KG, automatically generate queries, search the Web, and visualize the dependency graph.
@@ 2.3. Retrieval-Augmented Generation (RAG)
RAG (Lewis et al., 2020; Guu et al., 2020) augments the input questions for LLMs with information retrieved from documents, knowledge graphs, or vector databases, and is considered a useful method for addressing LLM hallucinations. As such, RAG provides extra knowledge sources for LLMs and greatly improves their performance in a variety of tasks, especially in the knowledge-intensive ones.
RAG applications leverage tools and function calls to perform their job. This feature is introduced by LLM multi-agent frameworks (i.e., LangChain (Chase, 2022) and Langroid (lan, 2024)) and recent versions of OpenAI models. [^2] https://platform.openai.com/docs/guides/text-generation/function-calling This feature enables users to provide function descriptions to the LLM. The LLM responds with a structured response, such as a JSON file containing the required arguments to invoke any available functions. These functions serve as action executors and provide the option to supply APIs that the LLM can query to obtain essential information for responding to users. For example, when a user asks about the current weather in a specific city and provides the LLM with a weather API call description, the LLM can augment its response with information retrieved from the API. This facilitates a seamless and efficient interaction between the user and the LLM and provides rich responses that avoid hallucinations.
@@ 3. Approach
DON’T forget to include the package name in your questions. After receiving this information, ensure to get the package version type.
THEN ask the user if they want to construct the dependency graph, and if so, use ”ConstructKGTool” to construct the dependency graph. Otherwise, say ”Could not retrieve package type or version” and NO_ANSWER.

### v2

@@ Abstract.
Managing software dependencies is a crucial maintenance task in software development. It is becoming a rapidly growing research field, especially in light of the significant increase in software supply chain attacks. Specialized expertise and substantial developer effort are required to fully comprehend dependencies and reveal hidden properties (e.g., number of dependencies, dependency chains, depth of dependencies). Recent advancements in Large Language Models (LLMs) allow the retrieval of information from various data sources for response generation, thus providing a new opportunity to uniquely manage software dependencies. To highlight the potential of this technology, we present DepsRAG, a proof-of-concept Retrieval Augmented Generation (RAG) approach that constructs direct and transitive dependencies of software packages as a Knowledge Graph (KG) in four popular software ecosystems. DepsRAG can answer user questions about software dependencies by automatically generating necessary queries to retrieve information from the KG, and then augmenting the input of LLMs with the retrieved information. DepsRAG can also perform a Web search to answer questions that the LLM cannot directly answer via the KG. We identify tangible benefits that DepsRAG can offer and discuss its limitations.  [^1] DepsRAG implementation available: https://github.com/Mohannadcse/DepsRAG.
copyright: noneconference: ; ; isbn:
@@ 1. Introduction
Software development increasingly relies on code reuse, which can be accomplished through software packages. Utilizing third-party packages to build software can significantly improve quality and productivity (Lim, 1994; Mohagheghi et al., 2004). These packages, dependencies, and their maintainers have formed large software ecosystems (Sonatype, 2021). Managing dependencies between packages is an emerging and important problem (Artho et al., 2012; Bogart et al., 2016; Decan et al., 2019), especially since open-source and third-party components have become widely adopted in software development. This new trend highlights the importance of software supply chain security, which is becoming more critical over time.
Today, developers are expected to manage their software dependencies (Kula et al., 2018). However, in a recent survey, developers reported that they are not confident in their current dependency management practices (Tidelift, [n. d.]). Although developers often have full control over which dependency packages and versions to use, they have no control over the dependencies of the packages used. Such transitive dependencies, which often amount to more than direct dependencies, may have various vulnerabilities that put software projects at risk. To mitigate these risks, we developed DepsRAG to support practitioners and software developers understand the dependencies in their software supply chain. DepsRAG facilitates constructing and analyzing software dependencies for four popular software ecosystems. It also answers key questions about the dependency graph, including the number of dependencies, the depth of the dependency graph, and the number of paths/chains. The use of LLM enables DepsRAG to autonomously generate queries for KGs, unlike prior methods demanding hardcoded or user-generated queries (Dann et al., 2023; Litzenberger et al., 2023; Yamaguchi et al., 2014), thereby enhancing scalability. This retrieved information enhances the user’s question, referred to as Retrieval Augmented Generation (RAG), enabling accurate LLM responses. DepsRAG works in two phases. In the first phase, the user provides the name and ecosystem of the project. Then it transitively constructs the dependencies as a KG. In the second phase, the user can chat with DepsRAG to answer questions about the dependencies. DepsRAG is equipped with several tools to facilitate its functionality to access the KG, automatically generate queries, search the Web, and visualize the dependency graph.
@@ 2.3. Retrieval-Augmented Generation (RAG)
RAG (Lewis et al., 2020; Guu et al., 2020) augments the input questions for LLMs with information retrieved from documents, knowledge graphs, or vector databases, and is considered a useful method for addressing LLM hallucinations. As such, RAG provides extra knowledge sources for LLMs and greatly improves their performance in a variety of tasks, especially in the knowledge-intensive ones. RAG applications leverage tools and function calls to perform their job. This feature is introduced by LLM multi-agent frameworks (i.e., LangChain (Chase, 2022) and Langroid (lan, 2024)) and recent versions of OpenAI models. [^2] https://platform.openai.com/docs/guides/text-generation/function-calling This feature enables users to provide function descriptions to the LLM. The LLM responds with a structured response, such as a JSON file containing the required arguments to invoke any available functions. These functions serve as action executors and provide the option to supply APIs that the LLM can query to obtain essential information for responding to users. For example, when a user asks about the current weather in a specific city and provides the LLM with a weather API call description, the LLM can augment its response with information retrieved from the API. This facilitates a seamless and efficient interaction between the user and the LLM and provides rich responses that avoid hallucinations.
@@ 3. Approach
DON’T forget to include the package name in your questions. After receiving this information, ensure to get the package version type. THEN ask the user if they want to construct the dependency graph, and if so, use ”ConstructKGTool” to construct the dependency graph. Otherwise, say ”Could not retrieve package type or version” and NO_ANSWER.